Given a test statistic of 1.09, you are asked to perform a hypothesis test at the α = 0:05 level of
significance for H0: p = 0.1 vs Ha: p =/= 0.1. Find the p-value for this test.

  • A.

    0.8622

  • B.

    0.7242

  • C.

    0.2758

  • D.

    0.1379

  • E. 1.7244
Answer #1

solution

this is two tailed test

P(z >1.09 ) = 1 - P(z <1.09 ) = 1-0.8621=0.1379*2=0.2758

P-value = 0.2758 ( using z table )